# Visual Whole-Body Control for Legged Loco-Manipulation

### v3

@@ V Conclusions, Limitations and Future Works
In this paper, we proposed a fully autonomous mobile manipulation system based on quadruped robots and a hierarchical training pipeline, Visual Whole-Body Control (VBC). VBC consists of a low-level goal-reaching policy and a high-level task-planning policy, which is trained through reinforcement learning (RL) and imitation learning (IL). Although we showed great success in picking various objects, we found several limitations of the system: 1) Compounding error. Our pipeline is hierarchical and requires each module to be coherent and precise to make the robot work well. 2) In-precise depth estimation. The RGBD camera does not reflect a perfect depth estimation, especially on reflective objects. 3) Bad gripper. The gripper of the Unitree Z1 Arm is a beak-like gripper instead of a parallel one, which tends to push objects away and is hard for accurate manipulation. 4) Unstable vision inputs. The tracking SAM that is used to provide the mask of annotated objects, may lose track due to camera deviation and occlusion, or be confused when the selected objects have a similar color to the other things. In our current pipeline, before the robot autonomously picks up objects, we must manually specify the target object in both cameras mounted on the robot. Thanks to the generalizability of trackingSAM, we find that as long as the object is visible to one camera (either the head or the wrist camera), the annotation mask can be propagated to another view for automated picking, and free the robot from being specified in the between of the range of the two cameras, which improves the usability of our system. Since the requirements of annotation and object visibility are totally decoupled from our sim-to-real framework, in the future we can further improve and replace these modules with automated annotation algorithms and a search/exploration module to further improve the automation of the whole system. Future works should also consider upgrading the hardware, like replacing a better gripper and RGBD cameras that provide better depth estimation; and simplify the working pipeline.
@@ II-B High-Level Training
The privileged teacher policy is trained through RL, specifically, we use PPO for training with designed reward functions. Reward functions We put all reward functions along with their weights of training the high-level policies in Tab. III. Except for $r_{\text{progress}}$, $r_{\text{completion}}$ and $r_{\text{completion}}$ that is determined by the task stage, we define a set of assistant rewards to help the algorithm converge faster. Among them, $r_{\text{acc}}$ limits the arm joints velocities’ change rate; $r_{\text{action}}$ helps smooth the output of the high-level policy to prevent it from changing quickly; $r_{\text{ee orn}}$ and $r_{\text{base orn}}$ encourages the arm and the robot body to be aimed at the object; $r_{\text{base approach}}$ encourages the robot to be close to the object when the robot is far away from the object. Beyond these rewards, we also design curriculums such as gradually adding some rewards, commanding the low-level agent using the action of the last step considering the delay on the real robot, and clipping the high-level action outputs to help safely real-world deployment.
@@
Additional Training Techniques We design several training techniques with our prior knowledge to help the high-level policy to learn a well-performed, stable, not only in the simulation, but can also be transferred into the real world. We briefly introduce them in the following:
@@
    Reward curriculum. The command penalty reward $r_{\text{cmd}}$ is designed to encourage the robot to slow its velocity when it is close to the object to be picked.
@@
    Randomly changes the object position. During training, we take a small probability ($10\%$) to randomly change the object position and pose. This helps the robot to learn the retrying behaviors when it finds it fails during one trial.

### v4

@@ V Conclusions, Limitations and Future Works
In this paper, we proposed a fully autonomous mobile manipulation system based on quadruped robots and a hierarchical training pipeline, Visual Whole-Body Control (VBC). VBC consists of a low-level goal-reaching policy and a high-level task-planning policy, which is trained through reinforcement learning and imitation learning. Although we showed great success in picking various objects, we found several failure cases due to the limitations of the system: 1) Compounding error. Our pipeline is hierarchical and requires each module to be coherent and precise to make the robot work well. 2) In-precise depth estimation. The RGBD camera does not reflect a perfect depth estimation, especially on reflective objects. 3) Bad gripper. The gripper of the Unitree Z1 Arm is a beak-like gripper instead of a parallel one, which tends to push objects away and is hard for accurate manipulation. 4) Unstable vision inputs. The tracking SAM that is used to provide the mask of annotated objects, may lose track due to camera deviation and occlusion, or be confused when the selected objects have a similar color to the other things. This is the most common cause leading to failure in our real-world experiments. In our current pipeline, before the robot autonomously picks up objects, we must manually specify the target object in both cameras mounted on the robot. Thanks to the generalizability of trackingSAM, we find that as long as the object is visible to one camera (either the head or the wrist camera), the annotation mask can be propagated to another view for automated picking, and free the robot from being specified in the between of the range of the two cameras, which improves the usability of our system. Since the requirements of annotation and object visibility are totally decoupled from our sim-to-real framework, in the future we can further improve and replace these modules with automated annotation algorithms and a search/exploration module to further improve the automation of the whole system. Future works should also consider upgrading the hardware, like replacing a better gripper and RGBD cameras that provide better depth estimation; and simplify the working pipeline.
@@ II-B High-Level Training
The privileged teacher policy is trained through RL, specifically, we use PPO for training with designed reward functions. Reward functions We put all reward functions along with their weights of training the high-level policies in Tab. III. Except for $r_{\text{progress}}$, $r_{\text{completion}}$ and $r_{\text{completion}}$ that is determined by the task stage, we define a set of assistant rewards to help the algorithm converge faster. Among them, $r_{\text{acc}}$ limits the arm joints velocities’ change rate; $r_{\text{cmd}}$ is designed to encourage the robot to slow its velocity when it is close to the object to be picked., $r_{\text{action}}$ helps smooth the output of the high-level policy to prevent it from changing quickly; $r_{\text{ee orn}}$ and $r_{\text{base orn}}$ encourages the arm and the robot body to be aimed at the object; $r_{\text{base approach}}$ encourages the robot to be close to the object when the robot is far away from the object.
@@
Additional Training Techniques We design several training techniques with our prior knowledge to help the high-level policy to be well-performed and stable, not only in the simulation, but can also be transferred into the real world. We briefly introduce them in the following:
@@
    Reward curriculum. The command penalty reward $r_{\text{cmd}}$ is added after the robot learns well to pick up objects.
@@
    Randomly changes the object position. During training, we take a small probability ($10\%$) to randomly change the object position and pose. This helps the robot to learn the retrying behaviors when it fails during one trial.